Directrix (directrix):

For what value of k are the roots of x^2 +kx + 600 = 0 consecutive negative integers?

OpenStudy (anonymous):

49

Directrix (directrix):

because ...?

OpenStudy (anonymous):

if u find out the divisors of 600,u can see only 24 and 25 are filling the condition

OpenStudy (anonymous):

so k will be -((-24)+(-25))

Directrix (directrix):

Yes, √600 = approx 24.5 so the consecutive integers must be -24 and -25. Or, x^2 + kx + 600 = (x+24)(x+25) = (x-24)(x-25). For consecutive negative integers, go with the first factoring. k = 49

OpenStudy (anonymous):

yeah

Directrix (directrix):

^5